# Rectified Diffusion Guidance for Conditional Generation

### v1

@@ Abstract
Classifier-Free Guidance (CFG), which combines the conditional and unconditional score functions with two coefficients summing to one, serves as a practical technique for diffusion model sampling. Theoretically, however, denoising with CFG cannot be expressed as a reciprocal diffusion process, which may consequently leave some hidden risks during use. In this work, we revisit the theory behind CFG and rigorously confirm that the improper configuration of the combination coefficients (i.e., the widely used summing-to-one version) brings about expectation shift of the generative distribution. To rectify this issue, we propose ReCFG  [^1] ReCFG, pronounced as “reconfigure”, is the abbreviation for “rectified Classifier-Free Guidance”. with a relaxation on the guidance coefficients such that denoising with ReCFG strictly aligns with the diffusion theory. We further show that our approach enjoys a closed-form solution given the guidance strength. That way, the rectified coefficients can be readily pre-computed via traversing the observed data, leaving the sampling speed barely affected. Empirical evidence on real-world data demonstrate the compatibility of our post-hoc design with existing state-of-the-art diffusion models, including both class-conditioned ones (e.g., EDM2 on ImageNet) and text-conditioned ones (e.g., SD3 on CC12M), without any retraining. We will open-source the code to facilitate further research.
@@ 1 Introduction
Diffusion probabilistic models (DPMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020), known simply as diffusion models, have achieved unprecedented capability improvement of high-resolution image generation. It is well recognized that, DPMs are the most prominent generative paradigm for a broad distribution (i.e., text-to-image generation) (Podell et al., 2024; Chen et al., 2024; Esser et al., 2024). Among DPM literature, Classifier-Free Guidance (CFG) (Ho & Salimans, 2021) serves as an essential factor, enabling better conditional sampling. Vanilla conditional sampling via DPMs introduces the conditional score function $s_{t}(\mathbf{x},c)=\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)$, resulting in poor performance in which synthesized samples appear to be visually incoherent and not faithful to the condition, even for large-scale models (Rombach et al., 2022). By drawing lessons from Bayesian theory, CFG employs an interpolation between conditional and unconditional score functions with a preset weight $\gamma$, i.e.,
$s_{t,\gamma}(\mathbf{x},c)=\gamma\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}),$ (1)
in which $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ is the unconditional score function by annihilating the condition effect. By doing so, DPMs turn out to formulate the underlying distribution with a gamma-powered distribution (Bradley & Nakkiran, 2024), i.e.,
@@
However, inspired by seminal works (Bradley & Nakkiran, 2024), we argue theoretically that denoising with CFG cannot be expressed as a reciprocal of vanilla diffusion process by adding Gaussian noises, since the expectation of score function of gamma-powered $q_{t,\gamma}(\mathbf{x}|c)$ is normally nonzero, violating the underlying theory of DPMs. Theoretically, score functions with zero expectation at all timesteps guarantee that the denoised $\tilde{\mathbf{x}}_{0}$ has expectation $\mathbb{E}[\tilde{\mathbf{x}}_{0}]=\frac{\alpha_{0}}{\alpha_{T}}\mathbb{E}[\mathbf{x}_{T}]$, thus $\mathbb{E}[\tilde{\mathbf{x}}_{0}]=\mathbb{E}[\mathbf{x}_{0}]$ and no bias on the conditional fidelity. Therefore, this theoretical flaw leaves some hidden risks during use, manifesting as a severe expectation shift phenomenon, i.e., the expectation of the gamma-powered distribution will be shifted away from the ground-truth of the conditional distribution $q_{t}(\mathbf{x}|c)$. This is more conspicuous when applying larger $\gamma$. Fig. 2 clearly clarifies the expectation shift, in which the peak of induced distribution via CFG in red fails to coincide with that of ground-truth $q_{0}(\mathbf{x}_{0}|c)$. This theoretical flaw is known in theory (Du et al., 2023; Karras et al., 2024a; Bradley & Nakkiran, 2024), while being largely ignored in practice.
Figure 1: Visualization of the lookup table on LDM (Rombach et al., 2022), EDM2 (Karras et al., 2024b), and SD3 (Esser et al., 2024), each of which consists of the expectation ratio $\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)]/\mathbb{E}_{\mathbf{x}_{t}}[\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)]$. Each pixel represents the scale of the pixel-wise ratio, i.e., color red implies that ratio is greater than one, while color blue stands for ratio smaller than one. The darker the color is, the farther the ratio appears away from one. The two images in one cell report the ratio of the first and the last denoising step under different NFEs.
@@
Aiming at consistency with the diffusion theory and thus better guidance efficacy, we specially design the constraints on $\gamma_{1}$ and $\gamma_{0}$, and theoretically confirm the feasibility. We further provide a closed-form solution to the constraints, and propose an algorithm to analytically determine $\gamma_{0}$ from a pre-computed lookup table in a post-hoc fashion. Thanks to the neat formulation, we can employ pixel-wise $\gamma_{0}$ according to the lookup table involving guidance strength $\gamma_{1}$, condition $c$ and timestep $t$, as demonstrated in Fig. 1. We name the above process ReCFG. Experiments with state-of-the-art DPMs, including both class-conditioned ones (e.g., EDM2 (Karras et al., 2024b)) and text-conditioned ones (e.g., SD3 (Esser et al., 2024)) under different NFEs and guidance strengths show that our ReCFG can achieve better guidance efficacy without retraining or extra time cost during inference stage. Hence, our work offers a new perspective on guided sampling of DPMs, encouraging more studies in the field of guided generation.
@@ 2 Related Work
DPMs and conditional generation. Diffusion probabilistic model (DPM) introduces a new scheme of generative modeling, formulated by forward diffusing and reverse denoising processes (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020). It is trained by optimizing the variational lower bound. Benefiting from this breakthrough, DPM achieves high generation fidelity, and even beat GANs on image generation. Conditional generation (Choi et al., 2021; Huang et al., 2023) takes better advantage of intrinsic intricate knowledge of data distribution, making DPM easier to scale up and the most promising option for generative modeling. Among the literature, text-to-image generation injects the embedding of text prompts to DPM, faithfully demonstrating the text content (Podell et al., 2024; Chen et al., 2024; Esser et al., 2024).
Classifier-Free Guidance. Classifier-Free Guidance (CFG) serves as the successor of Classifier Guidance (CG) (Dhariwal & Nichol, 2021), circumventing the usage of a classifier for noisy images. Both CFG and CG attempt to formulate the underlying distribution by concentrating more on condition influence, achieving better conditional fidelity. Despite great success in large-scale conditional generation, CFG faces a technical flaw that the guided distribution is not theoretically guaranteed to recover the ground-truth conditional distribution (Du et al., 2023; Karras et al., 2024a; Bradley & Nakkiran, 2024). There exists a shifting issue that the expectation of guided distribution is drifted away from the correct one. This phenomenon may harm the condition faithfulness, especially for extremely broad distribution (e.g., open-vocabulary synthesis).
@@ 3.1 Background on conditional DPMs and CFG
Let $\mathbf{x}_{0}\in\mathbb{R}^{D}$ be a $D$-dimensional random variable with an unknown distribution $q_{0}(\mathbf{x}_{0}|c)$, where $c\sim q(c)$ is the given condition. DPM (Sohl-Dickstein et al., 2015; Song et al., 2020; Ho et al., 2020) introduces a forward process $\{\mathbf{x}_{t}\}_{t\in(0,T]}$ by gradually corrupting data signal of $\mathbf{x}_{0}$ with Gaussian noise, i.e., the following transition distribution holds for any $t\in(0,T]$:
@@
in which $\alpha_{t},\sigma_{t}\in\mathbb{R}^{+}$ are differentiable functions of $t$ with bounded derivatives, referred to as the noise schedule. Let $q_{t}(\mathbf{x}_{t}|c)$ be the marginal distribution of $\mathbf{x}_{t}$ conditioned on $c$, DPM ensures that $q_{T}(\mathbf{x}_{T}|c)\approx\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ for some $\sigma>0$, and the signal-to-noise-ratio (SNR) $\alpha_{t}^{2}/\sigma_{t}^{2}$ is strictly decreasing with respect to timestep $t$ (Kingma et al., 2021).
Seminal works (Kingma et al., 2021; Song et al., 2020) studied the underlying stochastic differential equation (SDE) and ordinary differential equation (ODE) theory of DPM. The forward and reverse processes are as below for any $t\in[0,T]$:
$\mathrm{d}\mathbf{x}_{t}$ $=f(t)\mathbf{x}_{t}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}|c),$ (5)
$\mathrm{d}\mathbf{x}_{t}$ $=[f(t)\mathbf{x}_{t}-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)]\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}_{t},$ (6)
where $\mathbf{w}_{t},\bar{\mathbf{w}}_{t}$ are standard Wiener processes in forward and reverse time, respectively, and $f,g$ have closed-form expressions with respect to $\alpha_{t},\sigma_{t}$. The unknown $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)$ is referred to as the conditional score function. Probability flow ODE (PF-ODE) from Fokker-Planck equation enjoys the identical marginal distribution at each $t$ as that of the SDE in Eq. 6, i.e.,
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(t)\mathbf{x}_{t}-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c).$ (7)
@@
For better condition fidelity, during denoising stage, CFG (Ho & Salimans, 2021) turns to use a linear interpolation between conditional and unconditional score functions, i.e.,
$\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma}(\mathbf{x}_{t}|c)=\gamma\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)+(1-\gamma)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}).$ (9)
Then PF-ODE can be rewrote as
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f(t)\mathbf{x}_{t}-\frac{1}{2}g^{2}(t)\nabla_{\mathbf{x}_{t}}\log q_{t,\gamma}(\mathbf{x}_{t}|c).$ (10)
We further describe the CFG under the original DDIM theory. Recall that DDIM turns out to formulate non-Markovian forward diffusing process such that the reverse denoising process obeys the distribution with parameters $\{\delta_{t}\}_{t}$ (Song et al., 2021):
$q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)=q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})\sim\mathcal{N}\left(\alpha_{t-1}\mathbf{x}_{0}+\sqrt{\sigma_{t-1}^{2}-\delta_{t}^{2}}\cdot\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}},\delta_{t}^{2}\mathbf{I}\right).$ (11)
@@
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c)$ =$=\begin{cases}{q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta}^{t}(\mathbf{x}_{t},c),c)},&t>1,\\[5.0pt] {\mathcal{N}(\mathbf{f}_{\theta}^{t}(\mathbf{x}_{1}),\sigma_{1}^{2}\mathbf{I})},&t=1.\end{cases}$ (13)
DDIM proves that for any $\{\delta_{t}\}_{t}$, score matching of non-Markovian process above is equivalent to native DPM up to a constant. Under CFG setting with weight $\gamma$, we generalize the theory as below:
$\mathbf{f}_{\theta,\gamma}^{t}(\mathbf{x}_{t},c)$ $=\frac{1}{\alpha_{t}}(\mathbf{x}_{t}-\sigma_{t}(\gamma\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+(1-\gamma)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t))),$ (14)
$p_{\theta,\gamma}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c)$ =$=\begin{cases}{q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta,\gamma}^{t}(\mathbf{x}_{t},c),c)},&t>1,\\[5.0pt] {\mathcal{N}(\mathbf{f}_{\theta,\gamma}^{t}(\mathbf{x}_{1},c),\sigma_{1}^{2}\mathbf{I})},&t=1.\end{cases}$ (15)
@@
$J_{\delta,\gamma}(\bm{\epsilon}_{\theta})=\mathbb{E}_{q_{\delta}(\mathbf{x}_{0:T}|c)}[\log q_{\delta}(\mathbf{x}_{1:T}|\mathbf{x}_{0},c)-\log p_{\theta,\gamma}(\mathbf{x}_{0:T}|c)],$ (16)
@@
Figure 2: Visualization of expectation shift. The demonstrated toy data is simulated by $q_{0}(\mathbf{x}_{0}|c)\sim\mathcal{N}(c,1)$, $q(c)\sim\mathcal{N}(0,1)$, $q_{0}(\mathbf{x}_{0})\sim\mathcal{N}(0,2)$. Gamma-powered distribution $q_{0,\gamma}(\mathbf{x}_{0}|c)$ from CFG (Ho & Salimans, 2021) fails to recover the same conditional expectation as ground-truth due to expectation shift (i.e., probability density function and histogram by DDIM (Song et al., 2021) sampler in red). To make a further step, larger $\gamma$ suggests more severe expectation shift, i.e., the peak of $q_{0,\gamma}(\mathbf{x}_{0}|c)$ tends further away from $q_{0}(\mathbf{x}_{0}|c)$ (i.e., probability density function in blue) as $\gamma$ goes from $1.5$ to $2.5$. As a comparison, our ReCFG successfully recovers the ground-truth expectation and smaller variance (i.e., probability density function and histogram by DDIM (Song et al., 2021) sampler in green), consistent with the motivation of guided sampling.
3.2 Misconceptions about CFG on Expectation Shift
CFG is designed to concentrate on better exemplars for each denoising step by sharpening the gamma-powered distribution as below (Bradley & Nakkiran, 2024)

### v2

@@ Abstract
Classifier-Free Guidance (CFG), which combines the conditional and unconditional score functions with two coefficients summing to one, serves as a practical technique for diffusion model sampling. Theoretically, however, denoising with CFG cannot be expressed as a reciprocal diffusion process, which may consequently leave some hidden risks during use. In this work, we revisit the theory behind CFG and rigorously confirm that the improper configuration of the combination coefficients (i.e., the widely used summing-to-one version) brings about expectation shift of the generative distribution. To rectify this issue, we propose ReCFG  [^1] ReCFG, pronounced as “reconfigure”, is the abbreviation for “rectified Classifier-Free Guidance”. with a relaxation on the guidance coefficients such that denoising with ReCFG strictly aligns with the diffusion theory. We further show that our approach enjoys a closed-form solution given the guidance strength. That way, the rectified coefficients can be readily pre-computed via traversing the observed data, leaving the sampling speed barely affected. Empirical evidence on real-world data demonstrate the compatibility of our post-hoc design with existing state-of-the-art diffusion models, including both class-conditioned ones (e.g., EDM2 on ImageNet) and text-conditioned ones (e.g., SD3 on CC12M), without any retraining. Code is available at https://github.com/thuxmf/recfg.
@@ 1 Introduction
Diffusion probabilistic models (DPMs) [27, 13, 29], known simply as diffusion models, have achieved unprecedented capability improvement of high-resolution image generation. It is well recognized that, DPMs are the most prominent generative paradigm for a broad distribution (i.e., text-to-image generation) [22, 3, 9]. Among DPM literature, Classifier-Free Guidance (CFG) [12] serves as an essential factor, enabling better conditional sampling in various fields [25, 23]. Vanilla conditional sampling via DPMs introduces the conditional score function $s_{t}(\mathbf{x},c)=\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)$, resulting in poor performance in which synthesized samples appear to be visually incoherent and not faithful to the condition, even for large-scale models [25]. By drawing lessons from Bayesian theory, CFG employs an interpolation between conditional and unconditional score functions with a preset weight $\gamma$, i.e.,
$s_{t,\gamma}(\mathbf{x},c)=\gamma\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x}|c)+(1-\gamma)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x}),$ (1)
in which $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ is the unconditional score function by annihilating the condition effect. By doing so, DPMs turn out to formulate the underlying distribution with a gamma-powered distribution [1], i.e.,
@@
However, inspired by seminal works [1], we argue that denoising with CFG cannot be expressed as a reciprocal of vanilla diffusion process by adding Gaussian noises, since the normally nonzero score function expectation of gamma-powered $q_{t,\gamma}(\mathbf{x}|c)$ violates the underlying theory of DPMs. Theoretically, score functions with zero expectation at all timesteps guarantee that the denoised $\tilde{\mathbf{x}}_{0}$ has expectation $\mathbb{E}[\tilde{\mathbf{x}}_{0}]=\frac{\alpha_{0}}{\alpha_{T}}\mathbb{E}[\mathbf{x}_{T}]$, thus $\mathbb{E}[\tilde{\mathbf{x}}_{0}]=\mathbb{E}[\mathbf{x}_{0}]$ and no bias on the conditional fidelity. Therefore, this theoretical flaw leaves some hidden risks during use, manifesting as a severe expectation shift phenomenon, i.e., the expectation of the gamma-powered distribution will be shifted away from the ground-truth of the conditional distribution $q_{t}(\mathbf{x}|c)$. This is more conspicuous when applying larger $\gamma$. Fig. 1 clearly clarifies the expectation shift, in which the peak of induced distribution via CFG in red fails to coincide with that of ground-truth $q_{0}(\mathbf{x}_{0}|c)$. This theoretical flaw is known in theory [8, 15, 1], while being largely ignored in practice.
Figure 1: Visualization of expectation shift. The demonstrated toy data is simulated by $q_{0}(\mathbf{x}_{0}|c)\sim\mathcal{N}(c,1)$, $q(c)\sim\mathcal{N}(0,1)$, $q_{0}(\mathbf{x}_{0})\sim\mathcal{N}(0,2)$. Gamma-powered distribution $q_{0,\gamma}(\mathbf{x}_{0}|c)$ from CFG [12] fails to recover the same conditional expectation as ground-truth due to expectation shift (i.e., probability density function and histogram by DDIM [28] sampler in red). To make a further step, larger $\gamma$ suggests more severe expectation shift, i.e., the peak of $q_{0,\gamma}(\mathbf{x}_{0}|c)$ tends further away from $q_{0}(\mathbf{x}_{0}|c)$ (i.e., probability density function in blue) as $\gamma$ goes from $1.5$ to $2.5$. As a comparison, our ReCFG successfully recovers the ground-truth expectation and smaller variance (i.e., probability density function and histogram by DDIM [28] sampler in green), consistent with the motivation of guided sampling.
@@
Aiming at consistency with the diffusion theory and thus better guidance efficacy, we specially design the constraints on $\gamma_{1}$ and $\gamma_{0}$, and theoretically confirm the feasibility. We further provide a closed-form solution to the constraints, and propose an algorithm to analytically determine $\gamma_{0}$ from a pre-computed lookup table in a post-hoc fashion. Thanks to the neat formulation, we can employ pixel-wise $\gamma_{0}$ according to the lookup table involving guidance strength $\gamma_{1}$, condition $c$ and timestep $t$, as demonstrated in Fig. 2. We name the above process ReCFG. Compared with a global CFG weight applied on all denoising steps and pixels, ReCFG may achieve more flexible and accurate guidance. Experiments with state-of-the-art DPMs, including both class-conditioned ones (e.g., EDM2 [16]) and text-conditioned ones (e.g., SD3 [9]) under different NFEs and guidance strengths show that our ReCFG can achieve better guidance efficacy without retraining or extra time cost during inference stage. Hence, our work offers a new perspective on guided sampling of DPMs, encouraging more studies in the field of guided generation.
@@ 2 Related Work
DPMs and conditional generation. Diffusion probabilistic model (DPM) introduces a new scheme of generative modeling, formulated by forward diffusing and reverse denoising processes in a differential equation fashion [27, 13, 29]. Practically, it is trained by optimizing the variational lower bound. Benefiting from this breakthrough, DPM achieves high generation fidelity, and even beat GANs on image generation. By drawing lessons from conditional distribution, conditional generation [5, 14] takes better advantage of intrinsic intricate knowledge of data distribution, making DPM easier to scale up and the most promising option for generative modeling. Among the literature, text-to-image generation injects the embedding of text prompts to DPM, faithfully demonstrating the text content [22, 3, 9].
Classifier-Free Guidance. Classifier-Free Guidance (CFG) serves as the successor of Classifier Guidance (CG) [7], circumventing the usage of a classifier for noisy images. Both CFG and CG are based on Bayesian theory, and attempt to formulate the underlying distribution by concentrating more on condition influence, achieving better conditional fidelity. Despite great success in large-scale conditional generation, CFG faces a technical flaw that the guided distribution is not theoretically guaranteed to recover the ground-truth conditional distribution [8, 15, 1, 4]. To be more detailed, there exists a shifting issue that the expectation of guided distribution is drifted away from the correct one [1, 4]. This phenomenon may harm the condition faithfulness, especially for extremely broad distribution (e.g., open-vocabulary synthesis).
@@ 3.1 Background on conditional DPMs and CFG
Let $\mathbf{x}_{0}\in\mathbb{R}^{D}$ be a $D$-dimensional random variable with an unknown distribution $q_{0}(\mathbf{x}_{0}|c)$, where $c\sim q(c)$ is the given condition. DPM [27, 29, 13] introduces a forward process $\{\mathbf{x}_{t}\}_{t\in(0,T]}$ by gradually corrupting data signal of $\mathbf{x}_{0}$ with Gaussian noise, i.e., the following transition distribution holds for any $t\in(0,T]$:
@@
in which $\alpha_{t},\sigma_{t}\in\mathbb{R}^{+}$ are differentiable functions of $t$ with bounded derivatives, referred to as the noise schedule. Let $q_{t}(\mathbf{x}_{t}|c)$ be the marginal distribution of $\mathbf{x}_{t}$ conditioned on $c$, DPM ensures that $q_{T}(\mathbf{x}_{T}|c)\approx\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ for some $\sigma>0$, and the signal-to-noise-ratio (SNR) $\alpha_{t}^{2}/\sigma_{t}^{2}$ is strictly decreasing with respect to timestep $t$ [17].
Seminal works [17, 29] studied the underlying stochastic differential equation (SDE) and ordinary differential equation (ODE) theory of DPM. The forward and reverse processes are as below for any $t\in[0,T]$:
$\mathrm{d}\mathbf{x}_{t}$ $=f_{t}\mathbf{x}_{t}\mathrm{d}t+g_{t}\mathrm{d}\mathbf{w}_{t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}|c),$ (5)
$\mathrm{d}\mathbf{x}_{t}$ $=[f_{t}\mathbf{x}_{t}-g^{2}_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)]\mathrm{d}t+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t},$ (6)
where $\mathbf{w}_{t},\bar{\mathbf{w}}_{t}$ are standard Wiener processes in forward and reverse time, respectively, and $f_{t},g_{t}$ have closed-form expressions with respect to $\alpha_{t},\sigma_{t}$. The unknown $\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c)$ is referred to as the conditional score function. Probability flow ODE (PF-ODE) from Fokker-Planck equation enjoys the identical marginal distribution at each $t$ as that of the SDE in Eq. 6, i.e.,
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f_{t}\mathbf{x}_{t}-\frac{1}{2}g^{2}_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}|c).$ (7)
@@
For better condition fidelity, during denoising stage, CFG [12] turns to use a linear interpolation between conditional and unconditional score functions, i.e.,
$s_{t,\gamma}(\mathbf{x},c)=\gamma\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x}|c)+(1-\gamma)\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x}).$ (9)
Then PF-ODE can be rewritten as
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=f_{t}\mathbf{x}_{t}-\frac{1}{2}g^{2}_{t}s_{t,\gamma}(\mathbf{x}_{t},c).$ (10)
We further describe the CFG under the original DDIM theory. Recall that DDIM turns out to formulate discrete non-Markovian forward diffusing process such that the reverse denoising process obeys the distribution with parameters $\{\delta_{t}\}_{t=0}^{T}$ [28]:
$\;q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0},c)=q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ (11)
$\sim$ $\;\mathcal{N}\left(\alpha_{t-1}\mathbf{x}_{0}+\sqrt{\sigma_{t-1}^{2}-\delta_{t}^{2}}\cdot\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}},\delta_{t}^{2}\mathbf{I}\right).$ (12)
@@
$\hskip-5.69054ptp_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c)$ $=\begin{cases}{q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{f}_{\theta}^{t}(\mathbf{x}_{t},c),c)},&t>1,\\[2.0pt] {\mathcal{N}(\mathbf{f}_{\theta}^{t}(\mathbf{x}_{1}),\sigma_{1}^{2}\mathbf{I})},&t=1.\end{cases}$ (14)
DDIM proves that for any $\{\delta_{t}\}_{t}$, score matching of non-Markovian process above is equivalent to native DPM. With CFG weight $\gamma$, we generalize the theory as below:
$\hat{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},c,t)$ $=\gamma\bm{\epsilon}_{\theta}(\mathbf{x}_{t},c,t)+(1-\gamma)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (15)
$\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c)$ $=\frac{1}{\alpha_{t}}(\mathbf{x}_{t}-\sigma_{t}\hat{\bm{\epsilon}}_{\theta}(\mathbf{x}_{t},c,t)),$ (16)
$\hskip-5.69054pt\hat{p}_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},c)$ $=\begin{cases}{q_{\delta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{t},c),c)},&t>1,\\[2.0pt] {\mathcal{N}(\hat{\mathbf{f}}_{\theta}^{t}(\mathbf{x}_{1},c),\sigma_{1}^{2}\mathbf{I})},&t=1.\end{cases}$ (17)
@@
$J_{\delta,\gamma}(\bm{\epsilon}_{\theta})=\mathbb{E}_{q_{\delta}(\mathbf{x}_{0:T}|c)}\left[\log\frac{q_{\delta}(\mathbf{x}_{1:T}|\mathbf{x}_{0},c)}{\hat{p}_{\theta}(\mathbf{x}_{0:T}|c)}\right],$ (18)
@@
3.2 Misconceptions on Expectation Shift
CFG is designed to concentrate on better exemplars for each denoising step by sharpening the gamma-powered distribution as below [1]: